### v2

# Complexity results and active-set identification of a derivative-free method for bound-constrained problems

Abstract. In this paper, we analyze a derivative-free line search method designed for bound-constrained problems. Our analysis demonstrates that this method exhibits a worst-case complexity comparable to other derivative-free methods for unconstrained and linearly constrained problems. In particular, when minimizing a function with $n$ variables, we prove that at most $\operatorname{\mathcal{O}}{(n\epsilon^{-2})}$ iterations are needed to drive a criticality measure below a predefined threshold $\epsilon$, requiring at most $\operatorname{\mathcal{O}}{(n^{2}\epsilon^{-2})}$ function evaluations. We also show that the total number of iterations where the criticality measure is not below $\epsilon$ is upper bounded by $\operatorname{\mathcal{O}}{(n^{2}\epsilon^{-2})}$. Moreover, we investigate the method capability to identify active constraints at the final solutions. We show that, after a finite number of iterations, all the active constraints satisfying the strict complementarity condition are correctly identified.
Keywords. Worst-case complexity. Active-set identification. Derivative-free methods.
MSC2000 subject classifications. 90C56. 90C30. 90C26.
1 Introduction
Let us consider the following nonlinear bound-constrained optimization problem:
$\begin{split}&\min\ f(x)\\ &\ \ l_{i}\leq x_{i}\leq u_{i},\qquad i=1,\dots,n,\end{split}$ (1)
where $l_{i},u_{i}\in\mathbb{R}\cup\{\pm\infty\}$, $l_{i}<u_{i}$, $i=1,\dots,n$.
We restrict ourselves to considering problem (1) when function values are given by a time consuming black-box oracle. Hence, the analytical expression of $f$ is not available and first-order information cannot be explicitly used nor approximated within a reasonable amount of time. In such a context, derivative-free methods [2, 12, 25] are usually employed to solve the problem.
In the literature, several derivative-free methods have been proposed to solve problem (1) (even with more general constraints). In particular, we can distinguish among model-based methods [9, 12, 18, 20, 22, 33], where the objective function is sampled in a neighborhood of the current point to build an appropriate model to be minimized, direct-search methods [1, 19, 23, 24, 26], where the objective function is sampled in a neighborhood of the current point in order to find descent, and line search methods [28, 30], where directions are explored by allowing the stepsize to dinamically expand.
For model-based and direct-search methods applied to problems with linear constraints (thus including (1)), a worst-case analysis can be found in [19, 22], providing upper bounds on the maximum number of iterations and function evaluations needed to drive a criticality measure below a prespecified threshold. In particular, in [19], it is shown that at most $\operatorname{\mathcal{O}}{(n\epsilon^{-2})}$ iterations and $\operatorname{\mathcal{O}}{(n^{2}\epsilon^{-2})}$ function evaluations are needed, for a (deterministic) direct-search method, to produce the first point with a criticality measure below $\epsilon>0$, matching the same complexity for the unconstrained case [16, 34]. In [22], similar bounds of $\operatorname{\mathcal{O}}{(k_{D}^{2}\epsilon^{-2})}$ iterations and $\operatorname{\mathcal{O}}{(nk_{D}^{2}\epsilon^{-2})}$ function evaluations are obtained, matching the same complexity for the unconstrained case [17], with $k_{D}$ being a problem dimension-dependent constant which define a fully linear model.
In the current paper, we analyze a line search method to solve problem (1). The algorithm under analysis is a modification of the one proposed in [29], equipped with a line search technique described in [7]. In particular, for any considered direction (i.e., a vector of the canonical basis), the line search technique first checks for a sufficient decrease in the objective function using a giving stepsize. Then, if such a decrease is obtained, an extrapolation (or expansion) phase starts, where increasingly larger values of the stespize are tried until some conditions are met. This approach allows us to obtain complexity and identification results that extend those existing in the literature for direct-search [19] and model-based [22] methods. In particular, let us summarize the main contributions of this paper below.
  • The first contribution of the current paper is providing a worst-case analysis for the proposed line search method, which yields to the same bounds for direct search [19], that is, $\operatorname{\mathcal{O}}{(n\epsilon^{-2})}$ iterations and $\operatorname{\mathcal{O}}{(n^{2}\epsilon^{-2})}$ function evaluations to produce the first point with a criticality measure below $\epsilon>0$. Additionally, for the proposed algorithm, we are able to bound the total number of iterations where the criticality measure is not below $\epsilon$, thus going beyond the complexity results for direct-search methods given in [19].
  • The second contribution of the current paper is to show finite identification of the active constraints for the proposed line search method. Such a property is usually desirable for an optimization algorithm due to, among other things, the possibility of saving function evaluations if one recognizes the surfaces where a stationary points lies. Furthermore, in several applications, we might be interested only in the identification of the surface containing an optimal solution (or its support). In the literature, finite active-set identification was established for smooth optimization algorithms and proximal methods (see, e.g., [3, 5, 8, 21, 35]), also providing complexity bounds in some cases [6, 13, 32]. In a derivative-free setting, parameter-dependent estimates were used in [18, 27, 30], allowing for finite identification of active constraints if certain conditions hold. Moreover, finite identification results have been shown in [14] for a method using an inner approximation approach to minimize a function over the convex hull of a given set of vectors, meaning that, in finite time, the algorithm is able to identify, under appropriate assumptions, the vectors with zero weight in the convex combination representing the final solution.
    Here, we show that the proposed algorithm correctly identifies the active constraints satisfying the strict complementarity condition in a finite number of iterations, without using any parameter-dependent estimate. Namely, this feature is just an intrinsic property of the proposed algorithm. Such a result is obtained by using some tools from the analysis of derivative-based methods  [13, 32]. More specifically, we define a measure which represents the minimum strict complementarity among the active constraints, thereby providing a neighborhood of the limit points where the active-set identification holds. Let us remark that also this identification result is obtained thanks to the extrapolation techninque used in the line search procedure, which allows the stepsize to expand until we hit the border of the feasible set when we are in a neighborood of a stationary point.
1.1 Notations
Given $l,u\in\mathbb{R}^{n}$, we denote $[l,u]=\{x\in\mathbb{R}^{n}:\ l_{i}\leq x_{i}\leq u_{i},\ i=1,\dots,n\}$ the feasible set of problem (1). If ${\cal A}=\{a_{1},\dots,a_{p}\}\subset\mathbb{R}^{n}$ is a (finite) set of vectors, we denote
$\operatorname{cone}({\cal A})=\left\{x\in\mathbb{R}^{n}:\ x=\sum_{i=1}^{p}\beta_{i}a_{i},\ \beta_{i}\geq 0,\ i=1,\dots,p\right\}.$
Given $v\in\mathbb{R}^{n}$ and a set $S\subseteq\mathbb{R}^{n}$, we denote by $v_{S}$ the projection of $v$ onto $S$. Given $\beta\in\mathbb{R}$, we indicate the sign of $\beta$ by $\operatorname{sign}(\beta)$, that is, $\operatorname{sign}(\beta)$ is $-1$ if $\beta<0$, 00 if $\beta=0$ and $1$ if $\beta>0$. Finally, $\|v\|$ denotes the $\ell_{2}$-norm of vector $v$.
The paper is organized as follows. In Section 2, we define a line search algorithm for the solution of problem (1). Section 3 is devoted to the analysis of the asymptotic convergence of the proposed algorithm, followed by the derivation of worst-case complexity bounds. In Section 4, we show finite active-set identification of the proposed algorithm. Finally, in Section 5, we draw some conclusions.
2 The algorithm
This section is concerned with the definition of a line search algorithm to solve problem (1). The proposed method, denoted as Algorithm 1, is a slightly modified version of the one proposed in  [28] and uses some ideas from the one proposed in [7] for unconstrained problems.
At each iteration $k$, starting from the current iterate $x_{k}$, the algorithm sets $y_{k}^{1}=x_{k}$ and explores the coordinate directions ±$\pm e_{i}$, $i=1,\ldots,n$, using stepsizes $\nu_{k}^{i}=\max\{\tilde{\alpha}_{k}^{i},c\Delta_{k}\}$, where
$\Delta_{k}=\max_{i=1,\dots,n}\{\tilde{\alpha}_{k}^{i}\},$
and the quantities $\tilde{\alpha}_{k}^{i}$, $i=1,\dots,n$, are tentative stepsizes updated throughout the iterations. Then, the scheme produces actual stepsizes $\alpha_{k}^{i}$ to compute intermediate points $y_{k}^{i+1}=y_{k}^{i}+\alpha_{k}^{i}d_{k}^{i}$, with $d_{k}^{i}\in\{\pm e_{i}\}$, $i=1,\dots,n$. In particular, for any $y_{k}^{i}$, if $y_{k}^{i}+\nu_{k}^{i}d_{k}^{i}$ is infeasible (i.e., $y_{k}^{i}$ is too close to a bound) or does not provide a sufficient decrease in the objective function, then we set $\alpha_{k}^{i}=0$ (i.e., $y_{k}^{i+1}=y_{k}^{i}$). Otherwise, a sufficient decrease in the objective function is achieved by moving along $d_{k}^{i}$ with a feasible stepsize $\alpha_{k}^{i}$ determined by a line search procedure which will be described later. Then, we set $x_{k+1}=y_{k}^{n+1}$ and prepare for the next iteration $k+1$.
As a final note regarding our proposed scheme, we refer to $k$ as a successful iteration if $x_{k+1}\neq x_{k}$, indicating that at least one positive stepsize $\alpha_{k}^{i}$, $i=1,\ldots,n$, has been computed. Conversely, we refer to $k$ as an unsuccessful iteration if $x_{k+1}=x_{k}$, that is, if $\alpha_{k}^{i}=0$ for all $i=1,\ldots,n$. Depending on whether an iteration $k$ is successful or not, we use specific rules to update the tentative stepsizes for the next iteration $k+1$. In more detail, for a successful iteration $k$, each $\tilde{\alpha}_{k+1}^{i}$ is set to $\alpha_{k}^{i}$ if the latter is positive, whereas $\tilde{\alpha}_{k+1}^{i}$ is set to $\nu_{k}^{i}$ otherwise. For an unsuccessful iteration $k$, each $\tilde{\alpha}_{k+1}^{i}$ is set to $\theta\nu_{k}^{i}$, with $\theta\in(0,1)$.
Algorithm 1 Derivative-free line search algorithm
1:given $x_{0}\in[l,u]$, $\theta\in(0,1)$, $\delta\in(0,1)$, $\gamma>0$, $c\in(0,1]$, $\tilde{\alpha}_{0}^{i}>0$, $i=1,\dots,n$
2:for $k=0,1,\dots$ do
3:     set $\Delta_{k}=\max_{i=1,\dots,n}\{\tilde{\alpha}_{k}^{i}\}$
4:     set $y_{k}^{1}=x_{k}$
5:     for $i=1,\dots,n$ do
6:         set $\nu_{k}^{i}=\max\{\tilde{\alpha}_{k}^{i},c\Delta_{k}\}$
7:         compute $d_{k}^{i}$ and $\alpha_{k}^{i}$ by the line search$(y_{k}^{i},i,\gamma,\delta,\nu_{k}^{i})$
8:         set $y_{k}^{i+1}=y_{k}^{i}+\alpha_{k}^{i}d_{k}^{i}$
9:     end for
10:     set $x_{k+1}=y_{k}^{n+1}$
11:     if $x_{k+1}\neq x_{k}$ then
12:         set $\tilde{\alpha}_{k+1}^{i}=\begin{cases}\alpha_{k}^{i}\quad\text{if }\alpha_{k}^{i}>0\\ \nu_{k}^{i}\quad\text{otherwise}\end{cases}\quad i=1,\ldots,n$
13:     else
14:         set $\tilde{\alpha}_{k+1}^{i}=\theta\nu_{k}^{i}$, $i=1,\dots,n$
15:     end if
16:end for
Given a feasible point $x$, the exploration of the $i$th coordinate direction $e_{i}$ is performed by a line search procedure outlined in Algorithm 2. First, we check if the given stepsize $\nu$ is feasible along ±$\pm e_{i}$, that is, if one between $x+\nu e_{i}$ and $x-\nu e_{i}$ is feasible. If this is not the case, then we quit the line search returning a zero step length to indicate a failure. Otherwise, we try to determine if one between $e_{i}$ and $-e_{i}$ is a “good” descent direction, that is, if a sufficient decrease in the objective function can be obtained by using a feasible stepsize. If neither $e_{i}$ nor $-e_{i}$ qualifies as a suitable descent direction, then the line search procedure terminates, returning a zero step length to indicate a failure. Conversely, if a sufficient decrease of $f$ is obtained, then an extrapolation (or expansion) phase starts (i.e., lines 12–15), where we try to increase the stepsize to the maximum extent while preserving feasibility and guaranteeing the sufficient decrease condition. Specifically, the while loop keeps expanding the stepsize as long as the most recently accepted point remains strictly within the bounds (i.e., $\alpha<\alpha_{\text{max}}$) and the new tentative point is sufficiently better than the last accepted one (i.e., $f(x+\eta d)\leq f(x+\alpha d)-\gamma(\eta-\alpha)^{2}$).
Algorithm 2 line search($x$,$i$,$\gamma$,$\delta$,$\nu$)
1:if $\nu>\max\{u_{i}-x_{i},x_{i}-\ell_{i}\}$ then return $d=e_{i}$, $\alpha=0$
2:end if
3:set $\bar{\alpha}=\nu$
4:if $\bar{\alpha}\leq x_{i}-l_{i}$ and $f(x-\bar{\alpha}e_{i})\leq f(x)-\gamma\bar{\alpha}^{2}$ then
5:     set $d=-e_{i}$, $\alpha_{\text{max}}=x_{i}-l_{i}$ and go to line 12
6:end if
7:if $\bar{\alpha}\leq u_{i}-x_{i}$ and $f(x+\bar{\alpha}e_{i})\leq f(x)-\gamma\bar{\alpha}^{2}$ then
8:     set $d=e_{i}$, $\alpha_{\text{max}}=u_{i}-x_{i}$ and go to line 12
9:else
10:     return $d$ and $\alpha=0$
11:end if
12:set $\alpha=\bar{\alpha}$ and $\eta=\min\{\alpha/\delta,\alpha_{\text{max}}\}$
13:while ($\alpha<\alpha_{\text{max}}$ and $f(x+\eta d)\leq f(x+\alpha d)-\gamma(\eta-\alpha)^{2}$) do
14:     set $\alpha=\eta$ and $\eta=\min\{\alpha/\delta,\alpha_{\text{max}}\}$
15:end while
16:return $d$ and $\alpha$
3 Convergence and worst-case complexity
This section is devoted to the theoretical analysis of Algorithm 1.
3.1 Assumptions and preliminary results
Throughout the paper, the following assumptions will be always considered satisfied, even if not explicitly invoked.
Assumption 1.
The objective function $f\colon\mathbb{R}^{n}\to\mathbb{R}$ is continuously differentiable with a Lipschitz continuous gradient $\nabla f$ with constant $L>0$ over $[l,u]$, i.e.,
$\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|\quad\forall x,y\in[l,u].$
Moreover, $f$ is bounded from below over $[l,u]$, i.e., a constant $f_{\min}\in\mathbb{R}$ exists such that
$f_{\min}\leq f(x)\quad\forall x\in[l,u].$
Assumption 2.
A constant $M_{g}\geq 0$ exists such that
$\|\nabla f(x)\|\leq M_{g}$
for all $x\in[l,u]$.
Under Assumption 1, we can also provide the following definition of coordinate-wise Lipschitz constants.
Definition 1.
The coordinate-wise Lipschitz constants $L_{i}>0$, $i=1,\ldots,n$, of $\nabla f$ are such that, for all $x\in[l,u]$,
$\lvert\nabla_{i}f(x+se_{i})-\nabla_{i}f(x)\rvert\leq L_{i}\lvert s\rvert\quad\forall s\in\mathbb{R}\colon x+se_{i}\in[l,u],\quad i=1,\ldots,n.$
Moreover,
$L^{\text{max}}=\max_{i=1,\ldots,n}L_{i}.$ (2)
Now, given $x\in[l,u]$, let us introduce the following criticality measure:
$\chi(x)=\max_{\begin{subarray}{c}x+d\in[l,u]\\ \|d\|\leq 1\end{subarray}}-\nabla f(x)^{\top}d.$
The above measure has been successfully used in the analysis of some direct search methods for linearly constrained problems [19, 23, 24]. It can be interpreted as the progress on a first-order model in a ball centered at $x$ with unit ray subject to feasibility constraints [11, 24], thus generalizing $\|\nabla f(x)\|$ from the unconstrained setting. Originally proposed in [10] for more general constraints and further analyzed in [11], $\chi(x)$ is continuous, non-negative and such that $\chi(x)=0$ if and only if $x$ is a KKT point. So, we can define a stationary point as follows.
Definition 2.
A point $x^{*}\in[l,u]$ is said to be a stationary point of problem (1) if $\chi(x^{*})=0$.
Next, given $\epsilon\geq 0$, we define the set of $\epsilon$-active constraints at $x\in[l,u]$ as
$I_{l}(x,\epsilon)$ =$=\{i:x_{i}\leq l_{i}+\epsilon\},$
$I_{u}(x,\epsilon)$ =$=\{i:x_{i}\geq u_{i}-\epsilon\}.$
Namely, $I_{l}(x,\epsilon)$ and $I_{u}(x,\epsilon)$ denote the sets of lower and upper bound constraints, respectively, that are nearly active at $x$ with a tolerance $\epsilon$. Accordingly, let us define $N(x,\epsilon)$ as the $\epsilon$-normal cone generated by the $\epsilon$-active constraints, that is,
$N(x,\epsilon)=\operatorname{cone}\Big{(}{\{-e_{i},i\in I_{l}(x,\epsilon)\}\cup\{e_{i},i\in I_{u}(x,\epsilon)\}\cup\{0\}\Big{)}},$
while the $\epsilon$-tangent cone $T(x,\epsilon)$ is the polar of $N(x,\epsilon)$, that is,
$T(x,\epsilon)=N(x,\epsilon)^{\circ}=\{d\in\mathbb{R}^{n}:d^{\top}v\leq 0,\ \forall\ v\in N(x,\epsilon)\}.$
The use of $\epsilon$-normal and $\epsilon$-tangent cones is a well known tool in the analysis of direct search methods applied to linearly constrained problems [19, 23, 24]. Essentially, the set $x+T(x,\epsilon)$ is an approximation of the feasible region near a feasible point $x$, that is, moving from $x$ along any direction in $T(x,\epsilon)$ with a stepsize less than or equal to $\epsilon$ ensures that all constraints stay satisfied.
In our case, considering the structure of the feasible set of problem (1), it is straightforward to verify that a set of generators for $T(x,\epsilon)$ is given by
$G_{T(x,\epsilon)}=\{-e_{i},i\not\in I_{l}(x,\epsilon)\}\cup\{e_{i},i\not\in I_{u}(x,\epsilon)\}\cup\{0\},$ (3)
that is, $T(x,\epsilon)=\operatorname{cone}{(G_{T(x,\epsilon)})}$.
The following two propositions from [24] show how $\chi(x_{k})$ can be upper bounded by using of the projection of $-\nabla f(x_{k})$ onto $T(x_{k},\epsilon)$ and $N(x_{k},\epsilon)$.
Proposition 1 ([23, Proposition 8.2]).
If $x\in[l,u]$, then for all $\epsilon\geq 0$ we have that
$\chi(x)\leq\|(-\nabla f(x))_{T(x,\epsilon)}\|+\epsilon\sqrt{n}\|(-\nabla f(x))_{N(x,\epsilon)}\|.$
Proposition 2 ([23, Proposition 8.1]).
Given $\epsilon\geq 0$, let $G_{T(x,\epsilon)}$ be defined as in (3). If $(-\nabla f(x))_{T(x,\epsilon)}\neq 0$, then there exists $d\in G_{T(x,\epsilon)}$ such that
$\dfrac{1}{\sqrt{n}}\|(-\nabla f(x))_{T(x,\epsilon)}\|\leq-\nabla f(x)^{\top}d.$
In the convergence analysis of Algorithm 1, Propositions 12 will allow us to relate $\chi(x_{k})$ with $\Delta_{k+1}$ for every iteration $k$ (see Theorem 2 below). In particular, this will be obtained by applying the above results with $\epsilon=\Delta_{k}$ and using the following relation between $\Delta_{k}$ and $\Delta_{k+1}$.
Lemma 1.
Let $\{\Delta_{k}\}$ be the sequence of maximum tentative stepsizes produced by Algorithm 1. Then, for all $k$ we have
$\Delta_{k+1}\begin{cases}\geq\Delta_{k}\quad&\text{if }x_{k+1}\neq x_{k}\\ =\theta\Delta_{k}\quad&\text{if }x_{k+1}=x_{k}.\end{cases}$
Therefore, $\Delta_{k}\leq\Delta_{k+1}/\theta$ for all $k$.
Proof.
First, let us consider the case where $x_{k+1}\neq x_{k}$ (i.e., $k$ is a successful iteration). In this case, the algorithm sets $\tilde{\alpha}_{k+1}^{i}=\alpha_{k}^{i}$ if $\alpha_{k}^{i}>0$ and $\tilde{\alpha}_{k+1}^{i}=\nu_{k}^{i}$ if $\alpha_{k}^{i}=0$. Since, from the line search procedure, $\alpha_{k}^{i}\geq\nu_{k}^{i}$ for all $i=1,\ldots,n$, we can write
$\tilde{\alpha}_{k+1}^{i}\geq\nu_{k}^{i}\geq\tilde{\alpha}_{k}^{i},\quad i=1,\ldots,n,$
where the last inequality follows from the definition of $\nu_{k}^{i}$. Then, using the definition of $\Delta_{k}$, we get $\Delta_{k+1}\geq\Delta_{k}$.
Now, let us consider the case where $x_{k+1}=x_{k}$ (i.e., $k$ is an unsuccessful iteration). In this case, the algorithm sets $\tilde{\alpha}_{k+1}^{i}=\theta\nu_{k}^{i}=\theta\max\{\tilde{\alpha}_{k}^{i},c\Delta_{k}\}$. Namely, for all $i=1,\dots,n$, we have that
$\tilde{\alpha}_{k+1}^{i}=\begin{cases}\theta\tilde{\alpha}_{k}^{i}\quad&\text{if }\tilde{\alpha}_{k}^{i}\geq c\Delta_{k},\\ \theta c\Delta_{k}\quad&\text{if }\tilde{\alpha}_{k}^{i}<c\Delta_{k}.\end{cases}$ (4)
From the definition of $\Delta_{k}$ and the fact that $c\in(0,1]$, it follows that $\tilde{\alpha}_{k+1}^{i}\leq\theta\Delta_{k}$ for all $i=1,\ldots,n$, implying that
$\Delta_{k+1}=\max_{i=1,\ldots,n}{\tilde{\alpha}^{i}_{k+1}}\leq\theta\Delta_{k}.$ (5)
Now, let $\bar{\imath}\in\{1,\dots,n\}$ be such that $\tilde{\alpha}_{k}^{\bar{\imath}}=\Delta_{k}$. Since $c\in(0,1]$, we have $\tilde{\alpha}_{k}^{\bar{\imath}}\geq c\Delta_{k}$ and then, recalling (4), we get $\tilde{\alpha}_{k+1}^{\bar{\imath}}=\theta\tilde{\alpha}_{k}^{\bar{\imath}}=\theta\Delta_{k}$. It follows from (5) that
$\Delta_{k+1}=\tilde{\alpha}_{k+1}^{\bar{\imath}}=\theta\Delta_{k}.$
3.2 Global convergence
In the following proposition, we show that Algorithm 1 and Algorithm 2 are well defined, i.e., Algorithm 2 cannot cycle so that Algorithm 1 produces infinite sequences of points and stepsizes.
Proposition 3.
Algorithm 1 is well defined, i.e., it produces infinite sequences $\{x_{k}\}$, $\{\alpha_{k}^{i}\}$, $\{\tilde{\alpha}_{k}^{i}\}$, $i=1,\dots,n$.
Proof.
To prove that Algorithm 1 is well defined, we have to show that the line search procedure cannot infinitely cycle over steps 1315 in the while loop. Let us suppose, by contradiction, that the while loop does not terminate, i.e., we always have $\alpha<\alpha_{\text{max}}$. At the $j$-th iteration of the while loop, we have
$\alpha=\frac{\bar{\alpha}}{\delta^{j}}$
with $\delta<1$. If $\alpha_{\text{max}}$ is finite, then we have $\alpha>\alpha_{\text{max}}$, for $j$ sufficiently large, which is a contradiction. Otherwise, if $\alpha_{\text{max}}=+\infty$, then, for every $j$ we have
$f(x+\eta d)\leq f(x+\alpha d)-\gamma(\eta-\alpha)^{2},$ (6)
with $\eta=\alpha/\delta=\bar{\alpha}/\delta^{j+1}$. Now, for $j$ sufficiently large, (6) contradicts Assumption 1, i.e., that $f$ is bounded from below on the feasible set. ∎
In the following, we present some results concerning the global convergence of Algorithm 1 to stationary points. More specifically, by extending some results from [7, 28], we establish a relationship between $\nabla f(x_{k})$, with specific directions $d$, and the largest tentative step length at $\Delta_{k+1}$. As to be shown, the bound depends, besides on the problem dimension $n$, on the Lipschitz constant $L$ and the algorithm parameters $\gamma$, $\theta$ and $\delta$.
From now on, let us denote
$T_{k}:=T(x_{k},\Delta_{k})\quad\text{and}\quad G_{k}:=G(x_{k},\Delta_{k}),$ (7)
where $G_{k}$ is the set of generators of $T_{k}$ and is defined as in (3).
Theorem 1.
Let $\{x_{k}\}$ be the sequence produced by Algorithm 1. Then, for all $k$ and for all $d\in G_{k}$, we have that
$-\nabla f(x_{k})^{\top}d\leq\begin{cases}\left(\dfrac{\gamma+L}{\delta}+L\sqrt{n}\right)\Delta_{k+1}\quad&\text{if }x_{k+1}\neq x_{k},\\[11.99998pt] \left(\dfrac{\gamma+L^{\text{max}}}{\theta}\right)\Delta_{k+1}\quad&\text{if }x_{k+1}=x_{k}.\end{cases}$ (8)
Proof.
First, consider an iteration $k$ such that $x_{k+1}\neq x_{k}$ (i.e., a successful iteration). The following cases can occur, recalling that the analysis is limited to considering directions ±$\pm e_{i}$, $i=1,\ldots,n$, belonging to $G_{k}$.
  • (1a)
    $e_{i}\in G_{k}$, $y_{k}^{i+1}=l_{i}$ and $y_{k}^{i}=l_{i}$. Then, $\alpha_{k}^{i}=0$ and $\tilde{\alpha}_{k+1}^{i}=\nu_{k}^{i}$. From the instructions of the line search procedure, we have that
    $f(y_{k}^{i}+\tilde{\alpha}_{k+1}^{i}e_{i})>f(y_{k}^{i})-\gamma(\tilde{\alpha}_{k+1}^{i})^{2}.$
    By the mean value theorem, we have
    $f(y_{k}^{i}+\tilde{\alpha}_{k+1}^{i}e_{i})-f(y_{k}^{i})=\tilde{\alpha}_{k+1}^{i}\nabla_{i}f(\xi_{k}^{i}),$
    where $\xi_{k}^{i}=y_{k}^{i}+t_{k}^{i}\tilde{\alpha}_{k+1}^{i}e_{i}$ and $t_{k}^{i}\in(0,1)$. Then,
    $-\nabla_{i}f(\xi_{k}^{i})<\gamma{\tilde{\alpha}_{k+1}^{i}}.$
    It follows that
    $-\nabla_{i}f(\xi_{k}^{i})+\nabla_{i}f(x_{k})-\nabla_{i}f(x_{k})<\gamma{\tilde{\alpha}_{k+1}^{i}}$
    and we can write
    $\begin{split}-\nabla_{i}f(x_{k})&<\gamma{\tilde{\alpha}_{k+1}^{i}}-(\nabla_{i}f(x_{k})-\nabla_{i}f(\xi_{k}^{i}))\\ &\leq\gamma{\tilde{\alpha}_{k+1}^{i}}+L\|x_{k}-\xi_{k}^{i}\|\\ &\leq\gamma{\tilde{\alpha}_{k+1}^{i}}+L\|x_{k}-y_{k}^{i}\|+L\|y_{k}^{i}-\xi_{k}^{i}\|.\end{split}$ (9)
    Moreover, since $t_{k}^{i}\in(0,1)$, we have $t^{i}_{k}\tilde{\alpha}_{k+1}^{i}\leq{\tilde{\alpha}_{k+1}^{i}}$. Then,
    $\|y_{k}^{i}-\xi_{k}^{i}\|\leq{\tilde{\alpha}_{k+1}^{i}}.$
    Hence, since $\alpha_{k+1}^{i}\leq\tilde{\alpha}_{k+1}^{i}\leq\Delta_{k+1}$ for all $i=1,\dots,n$, and taking into account that $y_{k}^{i}=x_{k}+\sum_{j=1}^{i-1}\alpha_{k}^{j}d_{k}^{j}$, so that $\|x_{k}-y_{k}^{i}\|\leq\sqrt{n}\Delta_{k+1}$, we get
    $(-\nabla f(x_{k}))^{\top}e_{i}<\gamma{\tilde{\alpha}_{k+1}^{i}}+L\|x_{k}-y_{k}^{i}\|+L{\tilde{\alpha}_{k+1}^{i}}\leq\left({\gamma+L}+L\sqrt{n}\right)\Delta_{k+1}.$ (10)
  • (1b)
    $y_{k}^{i+1}=l_{i}$ and $y_{k}^{i}>l_{i}$. Then, $\tilde{\alpha}_{k+1}^{i}=\alpha_{k}^{i}>0$ and $-e_{i}\in G_{k}$. From the instructions of the line search procedure, there exists $\beta_{k}^{i}\in\{0,\delta\tilde{\alpha}_{k+1}^{i}\}$ such that
    $f(y_{k}^{i}-\tilde{\alpha}_{k+1}^{i}e_{i})\leq f(y_{k}^{i}-\beta_{k}^{i}e_{i})-\gamma(\tilde{\alpha}_{k+1}^{i}-\beta_{k}^{i})^{2}\leq f(y_{k}^{i}-\beta_{k}^{i}e_{i}).$
    By the mean value theorem, we have
    $f(y_{k}^{i}-\tilde{\alpha}_{k+1}^{i}e_{i})-f(y_{k}^{i}-\beta_{k}^{i}e_{i})=-(\tilde{\alpha}_{k+1}^{i}-\beta_{k}^{i})\nabla_{i}f(\xi_{k}^{i}),$
    where $\xi_{k}^{i}=y_{k}^{i}-\beta_{k}^{i}e_{i}+t_{k}^{i}(\beta_{k}^{i}-\tilde{\alpha}_{k+1}^{i})e_{i}$ and $t_{k}^{i}\in(0,1)$. Then,
    $-\nabla_{i}f(\xi_{k}^{i})\leq 0.$
    Moreover, from the definition of $\beta_{k}^{i}$ and the fact that $t^{i}_{k}\in(0,1)$, we have $(1-t^{i}_{k})\beta_{k}^{i}\leq(1-t_{k}^{i})\tilde{\alpha}_{k+1}^{i}$. Then,
    $\|y_{k}^{i}-\xi_{k}^{i}\|=(1-t^{i}_{k})\beta_{k}^{i}+t_{k}^{i}\tilde{\alpha}^{i}_{k+1}\leq\tilde{\alpha}^{i}_{k+1}.$
    Hence, by similar reasonings as in case (1a), we obtain
    $(-\nabla f(x_{k}))^{\top}(-e_{i})\leq\left(L+L\sqrt{n}\right)\Delta_{k+1}.$ (11)
  • (2a)
    $-e_{i}\in G_{k}$, $y_{k}^{i+1}=u_{i}$ and $y_{k}^{i}=u_{i}$. Then, $\alpha_{k}^{i}=0$ and $\tilde{\alpha}_{k+1}^{i}=\nu_{k}^{i}$. Reasoning as in case (1a), with minor differences, we get
    $(-\nabla f(x_{k}))^{\top}(-e_{i})\leq\left({\gamma+L}+L\sqrt{n}\right)\Delta_{k+1}.$ (12)
  • (2b)
    $y_{k}^{i+1}=u_{i}$ and $y_{k}^{i}<u_{i}$. Then, $\tilde{\alpha}_{k+1}^{i}=\alpha_{k}^{i}>0$ and $e_{i}\in G_{k}$. Reasoning as in case (1b), with minor differences, we get
    $(-\nabla f(x_{k}))^{\top}e_{i}\leq\left(L+L\sqrt{n}\right)\Delta_{k+1}.$ (13)
  • (3a)
    {±$\{\pm e_{i}\}\cap G_{k}\neq\emptyset$, $l_{i}<y_{k}^{i+1}<u_{i}$ and $y_{k}^{i}=y_{k}^{i+1}$. Then, $\alpha_{k}^{i}=0$ and $\tilde{\alpha}_{k+1}^{i}=\nu_{k}^{i}$. From the instructions of the line search procedure, we have that
    $\begin{split}f(y_{k}^{i}+\tilde{\alpha}_{k+1}^{i}e_{i})&>f(y_{k}^{i})-\gamma(\tilde{\alpha}_{k+1}^{i})^{2}\quad\text{if}\ e_{i}\in G_{k},\\ f(y_{k}^{i}-\tilde{\alpha}_{k+1}^{i}e_{i})&>f(y_{k}^{i})-\gamma(\tilde{\alpha}_{k+1}^{i})^{2}\quad\text{if}\ -e_{i}\in G_{k}.\end{split}$
    By the mean value theorem, we have
    $\begin{split}f(y_{k}^{i}+\tilde{\alpha}_{k+1}^{i}e_{i})-f(y_{k}^{i})&=\nabla_{i}f(\xi_{k}^{i})\tilde{\alpha}_{k+1}^{i}\quad\text{if}\ e_{i}\in G_{k},\\ f(y_{k}^{i}-\tilde{\alpha}_{k+1}^{i}e_{i})-f(y_{k}^{i})&=-\nabla_{i}f(\bar{\xi}_{k}^{i})\tilde{\alpha}_{k+1}^{i}\quad\text{if}\ -e_{i}\in G_{k}.\end{split}$
    where $\xi_{k}^{i}=y_{k}^{i}+t_{k}^{i}\tilde{\alpha}_{k+1}^{i}e_{i}$ and $\bar{\xi}_{k}^{i}=y_{k}^{i}-\bar{t}_{k}^{i}\tilde{\alpha}_{k+1}^{i}e_{i}$, with $t_{k}^{i},\bar{t}_{k}^{i}\in(0,1)$. Then,
    $\begin{split}-\nabla_{i}f(\xi_{k}^{i})&<\gamma\tilde{\alpha}_{k+1}^{i}\quad\text{if}\ e_{i}\in G_{k},\\ \nabla_{i}f(\bar{\xi}_{k}^{i})&<\gamma\tilde{\alpha}_{k+1}^{i}\quad\text{if}\ -e_{i}\in G_{k}.\end{split}$
    and, recalling that $t^{i}_{k},\bar{t}^{i}_{k}\in(0,1)$, we also have
    $\begin{split}\|y_{k}^{i}-\xi_{k}^{i}\|&=t_{k}^{i}\tilde{\alpha}_{k+1}^{i}\leq{\tilde{\alpha}_{k+1}^{i}}\quad\text{if}\ e_{i}\in G_{k},\\ \|y_{k}^{i}-\bar{\xi}_{k}^{i}\|&=t_{k}^{i}\tilde{\alpha}_{k+1}^{i}\leq{\tilde{\alpha}_{k+1}^{i}}\quad\text{if}\ -e_{i}\in G_{k}.\end{split}$
    Hence, reasoning as in case (1a) (i.e., using (9) for $e_{i}$ and applying minor changes to (9), with $\xi_{k}^{i}$ replaced by $\bar{\xi}_{k}^{i}$, for $-e_{i}$), we obtain
    $\begin{split}-\nabla f(x_{k})^{\top}e_{i}&\leq\left({\gamma+L}+L\sqrt{n}\right)\Delta_{k+1}\quad\text{if}\ e_{i}\in G_{k},\\ -\nabla f(x_{k})^{\top}(-e_{i})&\leq\left({\gamma+L}+L\sqrt{n}\right)\Delta_{k+1}\quad\text{if}\ -e_{i}\in G_{k}.\end{split}$ (14)
  • (3b)
    $l_{i}<y_{k}^{i+1}<u_{i}$ and $y_{k}^{i}\neq y_{k}^{i+1}$. Then, $\alpha_{k}^{i}>0$, $\tilde{\alpha}_{k+1}^{i}=\alpha_{k}^{i}$ and $d_{k}^{i}\in G_{k}$. Let us assume that $d_{k}^{i}=-e_{i}$, i.e., $y_{k}^{i+1}=y_{k}^{i}-\tilde{\alpha}_{k+1}^{i}e_{i}$ (the proof for the case $d_{k}^{i}=e_{i}$ is identical, except for minor changes). Then, from the instructions of the line search procedure, there exist $\beta_{k}^{i}$ and $\bar{\beta}_{k}^{i}$ such that $\beta_{k}^{i}\in\{0,\delta\tilde{\alpha}_{k+1}^{i}\}$, $\tilde{\alpha}_{k+1}^{i}<\bar{\beta}_{k}^{i}\leq\tilde{\alpha}_{k+1}^{i}/\delta$ and
    $\begin{split}f(y_{k}^{i}-\tilde{\alpha}_{k+1}^{i}e_{i})&\leq f(y_{k}^{i}-\beta_{k}^{i}e_{i})-\gamma(\tilde{\alpha}_{k+1}^{i}-\beta_{k}^{i})^{2}\leq f(y_{k}^{i}-\beta_{k}^{i}e_{i}),\\ f\left(y_{k}^{i}-\bar{\beta}_{k}^{i}e_{i}\right)&>f(y_{k}^{i}-\tilde{\alpha}_{k+1}^{i}e_{i})-\gamma\left(\bar{\beta}_{k}^{i}-\tilde{\alpha}_{k+1}^{i}\right)^{2}.\end{split}$
    By the mean value theorem, we have
    $\begin{split}f(y_{k}^{i}-\tilde{\alpha}_{k+1}^{i}e_{i})-f(y_{k}^{i}-\beta_{k}^{i}e_{i})&=-(\tilde{\alpha}_{k+1}^{i}-\beta_{k}^{i})\nabla_{i}f(\xi_{k}^{i}),\\ f\left(y_{k}^{i}-\bar{\beta}_{k}^{i}e_{i}\right)-f(y_{k}^{i}-\tilde{\alpha}_{k+1}^{i}e_{i})&=-(\bar{\beta}_{k}^{i}-\tilde{\alpha}_{k+1}^{i})\nabla_{i}f(\bar{\xi}_{k}^{i}),\end{split}$
    where $\xi_{k}^{i}=y_{k}^{i}-\beta_{k}^{i}e_{i}-t_{k}^{i}(\tilde{\alpha}_{k+1}^{i}-\beta_{k}^{i})e_{i}$ and $\bar{\xi}_{k}^{i}=y_{k}^{i}-\tilde{\alpha}_{k+1}^{i}e_{i}-\bar{t}_{k}^{i}(\bar{\beta}_{k}^{i}-\tilde{\alpha}_{k+1}^{i})e_{i}$, with $t_{k}^{i},\bar{t}_{k}^{i}\in(0,1)$. Then,
    $\begin{split}-\nabla_{i}f(\xi_{k}^{i})&\leq 0,\\ \nabla_{i}f(\bar{\xi}_{k}^{i})&<\gamma(\bar{\beta}_{k}^{i}-\tilde{\alpha}_{k+1}^{i}).\end{split}$
    Moreover, from the definition of $\bar{\beta}_{k}^{i}$, it follows that $0\leq\bar{\beta}_{k}^{i}-\tilde{\alpha}_{k+1}^{i}\leq(1/\delta-1)\tilde{\alpha}_{k+1}^{i}\leq\tilde{\alpha}_{k+1}^{i}/\delta$, where we have used the fact that $\delta\in(0,1)$. Then,
    $\begin{split}-\nabla_{i}f(\xi_{k}^{i})&\leq 0,\\ \nabla_{i}f(\bar{\xi}_{k}^{i})&<\gamma\left(\dfrac{1-\delta}{\delta}\right)\tilde{\alpha}_{k+1}^{i}\leq\gamma\frac{\tilde{\alpha}_{k+1}^{i}}{\delta},\end{split}$
    and, recalling that $t_{k}^{i},\bar{t}_{k}^{i}\in(0,1)$, we also have
    $\|y_{k}^{i}-\xi_{k}^{i}\|$ $=\beta_{k}^{i}+t_{k}^{i}(\tilde{\alpha}_{k+1}^{i}-\beta_{k}^{i})\leq\tilde{\alpha}_{k+1}^{i}\leq\frac{\tilde{\alpha}_{k+1}^{i}}{\delta},$
    $\|y_{k}^{i}-\bar{\xi}_{k}^{i}\|$ $=\tilde{\alpha}_{k+1}^{i}+t_{k}^{i}(\bar{\beta}_{k}^{i}-\tilde{\alpha}_{k+1}^{i})\leq\bar{\beta}_{k}^{i}\leq\frac{\tilde{\alpha}_{k+1}^{i}}{\delta}.$
    Hence, reasoning as in the previous case, we obtain
    $\begin{split}-\nabla f(x_{k})^{\top}(-e_{i})&=\nabla_{i}f(x_{k})\leq\left(\dfrac{\gamma+L}{\delta}+L\sqrt{n}\right)\Delta_{k+1},\\ -\nabla f(x_{k})^{\top}e_{i}&=-\nabla_{i}f(x_{k})\leq\left(\dfrac{\gamma+L}{\delta}+L\sqrt{n}\right)\Delta_{k+1}.\end{split}$ (15)
Hence, from (10), (11), (12), (13), (14) and (15), we conclude that
$-\nabla f(x_{k})^{\top}d\leq\left(\dfrac{\gamma+L}{\delta}+L\sqrt{n}\right)\Delta_{k+1}\quad\forall d\in G_{k}.$
Now, let us analyze an iteration $k$ such that $x_{k+1}=x_{k}$ (i.e., an unsuccessful iteration). In such a case, only cases (1a), (2a) and (3a) can occur, although we have to consider that $y_{k}^{i}=x_{k}$ and $\tilde{\alpha}_{k+1}^{i}=\theta\nu_{k}^{i}$. Hence, replacing $L$ with $L_{i}$, we have that the following relations hold:
$-\nabla f(x_{k})^{\top}e_{i}$ $\leq\frac{\gamma+L^{\text{max}}}{\theta}\Delta_{k+1}\quad\text{if}\ e_{i}\in G_{k},$
$-\nabla f(x_{k})^{\top}(-e_{i})$ $\leq\frac{\gamma+L^{\text{max}}}{\theta}\Delta_{k+1}\quad\text{if}\ -e_{i}\in G_{k}.$
As above, we finally conclude
$-\nabla f(x_{k})^{\top}d\leq\frac{\gamma+L^{\text{max}}}{\theta}\Delta_{k+1}\quad\forall d\in G_{k}.$
Remark 1.
The result expressed in Theorem 1 strongly relies on the extrapolation phase of the line search procedure (i.e., lines 12–15 of Algorithm 2). In particular, since we quit the expansion with a failure in the objective decrease when we do not hit the border of the feasible set, then we are able to upper bound $\chi(x_{k})$ for all iterations, including the successful ones. This represents a relevant difference over direct-search [19] methods, where $\chi(x_{k})$ is usually upper bounded only for the unsuccessful iterations. Moreover, this property will allow us to give a complexity bound on the total number of iterations where $\chi(x_{k})$ is above a prespecified threshold (see Theorem 5 below).
Combining Proposition 2 and Theorem 1, it is now straightforward to relate $\chi(x_{k})$ with $\Delta_{k+1}$, as stated in the following result.
Theorem 2.
Let $\{x_{k}\}$ be the sequence of points produced by Algorithm 1. Then,
$\chi(x_{k})\leq\begin{cases}\sqrt{n}\left(\dfrac{\gamma+L}{\delta}+L\sqrt{n}+\dfrac{M_{g}}{\theta}\right)\Delta_{k+1}\quad&\text{if }x_{k+1}\neq x_{k},\\[11.99998pt] \sqrt{n}\left(\dfrac{\gamma+L^{\text{max}}+M_{g}}{\theta}\right)\Delta_{k+1}\quad&\text{if }x_{k+1}=x_{k}.\end{cases}$
Proof.
Using Proposition 1 with $\epsilon=\Delta_{k}$ and Lemma 1, we obtain
$\chi(x_{k})\leq\|(-\nabla f(x_{k}))_{T_{k}}\|+\frac{\Delta_{k+1}}{\theta}\sqrt{n}M_{g},$
where we have used the fact that $\|(-\nabla f(x_{k}))_{N(x_{k},\epsilon)}\|\leq\|\nabla f(x_{k})\|\leq M_{g}$, where the last inequality follows from Assumption 2. So, using Proposition 2, it follows that there exists a direction $d\in G_{k}$ such that
$\chi(x_{k})\leq\sqrt{n}\left(-\nabla f(x_{k})^{\top}d+\frac{\Delta_{k+1}}{\theta}\right)M_{g}.$
The desired result hence follows from Theorem 1. ∎
In order to get convergence to stationary points and provide worst-case complexity bounds for the proposed algorithm, for each iteration $k$ let us define
$\Phi_{k}=f(x_{k})+\frac{1}{2}c^{2}\gamma\Delta_{k}^{2}.$ (16)
Recalling Assumption 1, note that
$\Phi_{k}\geq f_{\min}\quad\forall k\geq 0.$ (17)
Now, in the next theorem, we bound the difference $\Phi_{k}-\Phi_{k-1}$ for each iteration $k$.
Theorem 3.
Let $\{x_{k}\}$ be the sequence produced by Algorithm 1. Then,
$\Phi_{k}-\Phi_{k-1}\leq-c_{1}\Delta_{k}^{2}\quad\forall k\geq 1,$ (18)
with
$c_{1}=\gamma\min\left\{\left(1-\frac{c^{2}}{2}\right),\frac{c^{2}}{2},\frac{1}{2}c^{2}\left(\frac{1-\theta^{2}}{\theta^{2}}\right)\right\}>0,$ (19)
where $c$ and $\theta$ are defined in Algorithm 1.
Proof.
For each iteration $k\geq 1$, consider the iteration $k-1$. The following cases can occur.
  • $x_{k}\neq x_{k-1}$ (i.e., $k-1$ is a successful iteration) and $\Delta_{k}=\Delta_{k-1}$. Hence, there exists a coordinate ${\bar{\imath}}$ such that we have moved from $x_{k-1}$ along ±$\pm e_{\bar{\imath}}$ and we have
    $\alpha_{k-1}^{\bar{\imath}}\geq\nu_{k-1}^{\bar{\imath}}=\max\left\{\tilde{\alpha}_{k-1}^{\bar{\imath}},c\Delta_{k-1}\right\}\geq c\Delta_{k-1}=c\Delta_{k}.$
    Then, from the line search procedure, we can write
    $f(x_{k})\leq f(x_{k-1})-\gamma c^{2}\Delta_{k}^{2}.$
    Hence, we have
    $\begin{split}\Phi_{k}-\Phi_{k-1}&=f(x_{k})-f(x_{k-1})+\frac{1}{2}c^{2}\gamma\left(\Delta_{k}^{2}-\Delta_{k-1}^{2}\right)\\ &\leq-\gamma c^{2}\Delta_{k}^{2}+\frac{1}{2}c^{2}\gamma\left(\Delta_{k}^{2}-\Delta_{k-1}^{2}\right)\\ &\leq-\gamma c^{2}\Delta_{k}^{2}+\frac{1}{2}c^{2}\gamma\Delta_{k}^{2}\\ &=-\gamma\frac{c^{2}}{2}\Delta_{k}^{2}.\end{split}$ (20)
  • $x_{k}\neq x_{k-1}$ (i.e., $k-1$ is a successful iteration) and $\Delta_{k}>\Delta_{k-1}$. Hence, there exists a coordinate ${\bar{\imath}}$ such that
    $\alpha_{k-1}^{\bar{\imath}}=\Delta_{k}.$
    Then, from the line search procedure, we can write
    $f(x_{k})\leq f(x_{k-1})-\gamma\Delta_{k}^{2}.$
    Hence, we have
    $\begin{split}\Phi_{k}-\Phi_{k-1}&=f(x_{k})-f(x_{k-1})+\frac{1}{2}c^{2}\gamma\left(\Delta_{k}^{2}-\Delta_{k-1}^{2}\right)\\ &\leq-\gamma\Delta_{k}^{2}+\frac{1}{2}c^{2}\gamma\left(\Delta_{k}^{2}-\Delta_{k-1}^{2}\right)\\ &\leq-\gamma\Delta_{k}^{2}+\frac{1}{2}c^{2}\gamma\Delta_{k}^{2}\\ &=-\gamma\left(1-\frac{c^{2}}{2}\right)\Delta_{k}^{2}.\end{split}$ (21)
  • $x_{k}=x_{k-1}$ (i.e., $k-1$ is an unsuccessful iteration). Hence, for all $i\in\{1,\dots,n\}$ we have either
    $\tilde{\alpha}_{k}^{i}=\theta\tilde{\alpha}_{k-1}^{i}\leq\theta\Delta_{k-1}$
    or
    $\tilde{\alpha}_{k}^{i}=\theta c\Delta_{k-1}\leq\theta\Delta_{k-1}$
    since $c\in(0,1]$. Hence, by Lemma 1, we have
    $\Delta_{k}=\theta\Delta_{k-1}.$
    Then,
    $\begin{split}\Phi_{k}-\Phi_{k-1}&=f(x_{k})-f(x_{k-1})+\frac{1}{2}c^{2}\gamma\left(\Delta_{k}^{2}-\Delta_{k-1}^{2}\right)\\ &\leq\frac{1}{2}c^{2}\gamma\left(\Delta_{k}^{2}-\frac{1}{\theta^{2}}\Delta_{k}^{2}\right)\\ &=-\frac{1}{2}c^{2}\gamma\left(\frac{1-\theta^{2}}{\theta^{2}}\right)\Delta_{k}^{2}.\end{split}$ (22)
Finally, we get (18) by combining (20), (21) and (22), where $c_{1}>0$ since $c\in(0,1]$. ∎
Using the above result, we can easily show the convergence to zero of the sequences of tentative and actual stepsizes produced by the algorithm, i.e., $\{\tilde{\alpha}_{k}^{i}\}$, $\{\alpha_{k}^{i}\}$, $i=1,\dots,n$, respectively, together with the convergence to zero of the sequence of maximum tentantive stepsizes $\{\Delta_{k}\}$.
Proposition 4.
Let $\{\tilde{\alpha}_{k}^{i}\}$, $\{\alpha_{k}^{i}\}$, $i=1,\dots,n$, and $\{\Delta_{k}\}$ be the sequences produced by Algorithm 1. We have that
$\lim_{k\to\infty}\tilde{\alpha}^{i}_{k}=0,$ $\quad i=1,\ldots,n;$ (23)
$\lim_{k\to\infty}\alpha_{k}^{i}=0,$ $\quad i=1,\ldots,n;$ (24)
$\lim_{k\to\infty}\Delta_{k}=0.$ (25)
Proof.
From (17) and (18), we get (25). Then, using the definition of $\Delta_{k}$ given in Algorithm 1, also (23) holds. Since, for all $k\geq 1$ and for all $i=1,\ldots,n$, from the instructions of the algorithm either $\alpha^{i}_{k-1}=0$ or $\alpha^{i}_{k-1}=\tilde{\alpha}^{i}_{k}$, then (24) follows from (23). ∎
Using Theorem 2 and Proposition 4, it is now possible to prove the convergence of the algorithm to stationary points.
Theorem 4.
Let $\{x_{k}\}$ be the sequence of points produced by Algorithm 1. Then,
  • $\lim_{k\to\infty}\chi(x_{k})=0$, i.e., every limit point of $\{x_{k}\}$ is stationary,
  • ${\lim_{k\to\infty}\|x_{k+1}-x_{k}\|=0}$.
Proof.
From Theorem 2 and (25) in Proposition 4, taking the limit for $k\to\infty$ it follows that $\chi(x_{k})\to 0$, that is, every limit point of $\{x_{k}\}$ is stationary. Finally, since $x_{k+1}-x_{k}=\sum_{i=1}^{n}\alpha_{k}^{i}d_{k}^{i}$, using (24) in Proposition 4 we also have that $\|x_{k+1}-x_{k}\|\to 0$. ∎
3.3 Worst-case complexity
This section is devoted to analyze the worst-case complexity of Algorithm 1. In particular,
  • (i)
    we give an upper bound of $\operatorname{\mathcal{O}}{(n^{2}\epsilon^{-2})}$ on the total number of iterations where $\chi(x_{k})$ is not below a prespecified threshold $\epsilon$;
  • (ii)
    we give an upper bound of $\operatorname{\mathcal{O}}{(n\epsilon^{-2})}$ on the number of iterations required to generate the first point $x_{k}$ where $\chi(x_{k})$ is below a prespecified tolerance $\epsilon$;
  • (iii)
    we give an upper bound of $\operatorname{\mathcal{O}}{(n^{2}\epsilon^{-2})}$ on the number of function evaluations required to generate the first point $x_{k}$ where $\chi(x_{k})$ is below a prespecified tolerance $\epsilon$.
We start by providing an upper bound of $\operatorname{\mathcal{O}}{(n^{2}\epsilon^{-2})}$ on the total number of iterations where $\chi(x_{k})\geq\epsilon$, with a given $\epsilon>0$.
Theorem 5.
Let $\{x_{k}\}$ be the sequence of points produced by the algorithm. Given any $\epsilon>0$, let
$K_{\epsilon}=\{k\colon\chi(x_{k})\geq\epsilon\}.$
Then, $|K_{\epsilon}|\leq\operatorname{\mathcal{O}}{(n^{2}\epsilon^{-2})}$. In particular,
$|K_{\epsilon}|\leq\left\lfloor\frac{c_{2}^{2}(\Phi_{0}-f_{\min})}{c_{1}}\epsilon^{-2}\right\rfloor,$
where $c_{1}$ is defined as in Theorem 3 and
$c_{2}=\sqrt{n}\max\left\{\frac{\gamma+L}{\delta}+L\sqrt{n}+\frac{M_{g}}{\theta},\dfrac{\gamma+L^{\text{max}}+M_{g}}{\theta}\right\}.$
Proof.
From Theorem 3, it follows that the sequence $\{\Phi_{k}\}$ is monotonically non-increasing. Furthermore, for all $k\geq 1$, we can write
$\Phi_{k}-\Phi_{0}\leq-c_{1}\sum_{j=1}^{k}\Delta_{j}^{2}=-c_{1}\sum_{j=0}^{k}\Delta_{j+1}^{2}.$ (26)
Since the sequence $\{\Phi_{k}\}$ is bounded from below, there exists $\Phi^{*}$ such that
$\lim_{k\to\infty}\Phi_{k}=\Phi^{*}\geq f_{\min},$
with $f_{\min}$ defined as in (17). Taking the limit for $k\to\infty$ in (26) we obtain
$\Phi_{0}-f_{\min}\geq\Phi_{0}-\Phi^{*}\geq c_{1}\sum_{k=0}^{\infty}\Delta_{k+1}^{2}\geq c_{1}\sum_{k\in K_{\epsilon}}\Delta_{k+1}^{2}.$
Therefore, using the definition of $K_{\epsilon}$ and Theorem 2, we get
$\Phi_{0}-f_{\min}\geq c_{1}\sum_{{k\in K_{\epsilon}}}\Delta_{k+1}^{2}\geq|K_{\epsilon}|c_{1}\frac{\epsilon^{2}}{c_{2}^{2}}.$
Thus, the desired result is obtained. ∎
As appears from the proof of Theorem 5, the above result relies on Theorem 2 which, in turn, uses Theorem 8. The latter, as pointed out in Remark 1, strongly relies on the extrapolation phase of the line search procedure (i.e., lines 12–15 of Algorithm 2), which allows us to bound $\chi(x_{k})$ at both successful and unsuccessful iterations.
In the following theorem, we give an upper bound of $\operatorname{\mathcal{O}}{(n\epsilon^{-2})}$ on the maximum number of iterations required to produce a point $x_{k}$ such that $\chi(x_{k})$ is below a given threshold $\epsilon>0$. This bound aligns with established findings for direct-search [19] and model-based [22] methods.
Theorem 6.
Let $\{x_{k}\}$ be the sequence of points produced by Algorithm 1. Given any $\epsilon>0$, let $j_{\epsilon}\geq 1$ be the first iteration such that $\chi(x_{j_{\epsilon}})<\epsilon$, that is, $\chi(x_{k})\geq\epsilon$ for all $k\in\{0,\dots,j_{\epsilon}-1\}$.
Then, $j_{\epsilon}\leq\operatorname{\mathcal{O}}{(n\epsilon^{-2})}$. In particular,
$j_{\epsilon}\leq\biggl{\lfloor}\frac{nc_{3}^{2}(\Phi_{0}-f_{\min})}{c_{1}}\epsilon^{-2}\biggr{\rfloor},$
where $c_{1}$ is given by (19) and
$c_{3}=\frac{\gamma+L^{\text{max}}+M_{g}}{\theta}.$ (27)
Proof.
Let $\Phi_{k}$ the function defined in (16). We can write
$\Phi_{j_{\epsilon}}-\Phi_{0}=\sum_{k=0}^{j_{\epsilon}-1}(\Phi_{k+1}-\Phi_{k})$
and, using Theorem 3, we have that
$\Phi_{j_{\epsilon}}-\Phi_{0}\leq-c_{1}\sum_{k=0}^{j_{\epsilon}-1}\Delta_{k+1}^{2}.$
Recalling (17) and the fact that $\Phi_{k}\geq f(x_{k})$ for $k\geq 0$, we get
$f_{\min}-\Phi_{0}\leq\Phi_{j_{\epsilon}}-\Phi_{0}\leq-c_{1}\sum_{k=0}^{j_{\epsilon}-1}\Delta_{k+1}^{2}.$ (28)
Now, we can partition the set of iteration indices $\{0,\dots,j_{\epsilon}-1\}$ into $\mathcal{S}_{j_{\epsilon}}$ and $\mathcal{U}_{j_{\epsilon}}$ such that
$k\in\mathcal{S}_{j_{\epsilon}}\,\Leftrightarrow\,x_{k}\not=x_{k-1},\quad k\in\mathcal{U}_{j_{\epsilon}}\,\Leftrightarrow\,x_{k}=x_{k-1},\quad\mathcal{S}_{j_{\epsilon}}\cup\mathcal{U}_{j_{\epsilon}}=\{0,\dots,j_{\epsilon}-1\},$
that is, $\mathcal{S}_{j_{\epsilon}}$ and $\mathcal{U}_{j_{\epsilon}}$ contain the successful and unsuccessful iterations up to $j_{\epsilon}-1$, respectively. So, from (28), we can write
$\Phi_{0}-f_{\min}\geq c_{1}\sum_{k\in\mathcal{S}_{j_{\epsilon}}}\Delta_{k+1}^{2}+c_{1}\sum_{k\in\mathcal{U}_{j_{\epsilon}}}\Delta_{k+1}^{2}.$ (29)
For all $k\in\mathcal{S}_{j_{\epsilon}}$, let us define the index $m(k)$ as follows:
  • if $\mathcal{U}_{j_{\epsilon}}\cap\{0,\ldots,k-1\}\not=\emptyset$, then $m(k)$ in the largest index of $\mathcal{U}_{j_{\epsilon}}\cap\{0,\ldots,k-1\}$;
  • otherwise, $m(k)=-1$.
Note that, by definition, $m(k)$ is the last unsuccessful iteration before iteration $k$, i.e., all the iterations from $m(k)+1$ to $k$ are successful iterations. Lemma 1 guarantees that $\Delta_{k+1}\geq\Delta_{m(k)+1}$ for all $k\in\mathcal{S}_{j_{\epsilon}}$. Using (29), we obtain
$\Phi_{0}-f_{\min}\geq c_{1}\sum_{k\in\mathcal{S}_{j_{\epsilon}}}\Delta_{m(k)+1}^{2}+c_{1}\sum_{k\in\mathcal{U}_{j_{\epsilon}}}\Delta_{k+1}^{2}.$
From Theorem 2, we have that $\Delta_{m(k)+1}\geq\chi(x_{m(k)})/(\sqrt{n}c_{3})$ for all $k\in\mathcal{S}_{j_{\epsilon}}$ and $\Delta_{k+1}\geq\chi(x_{k})/(\sqrt{n}c_{2})$ for all $k\in\mathcal{U}_{j_{\epsilon}}$. Since $\chi(x_{k})\geq\epsilon$ for all $k\in\{0,\dots,j_{\epsilon}-1\}$, with $\mathcal{S}_{j_{\epsilon}}\cup\mathcal{U}_{j_{\epsilon}}=\{0,\dots,j_{\epsilon}-1\}$, we get
$\Phi_{0}-f_{\min}\geq j_{\epsilon}\frac{c_{1}}{nc_{3}^{2}}\epsilon^{2}.$
Thus, the desired result is obtained. ∎
The last complexity result we give is about the maximum number of function evaluations required to produce a point $x_{k}$ such that $\chi(x_{k})$ is less than or equal to a given threshold $\epsilon>0$. Using arguments from the related literature [7, 34], we obtain an upper bound of $\operatorname{\mathcal{O}}{(n^{2}\epsilon^{-2})}$, which still aligns with established findings for direct-search [19] and model-based [22] methods
Theorem 7.
Let $\{x_{k}\}$ be the sequence of points produced by Algorithm 1. Given any $\epsilon>0$, let $j_{\epsilon}\geq 1$ be the first iteration such that $\chi(x_{j_{\epsilon}})<\epsilon$, that is, $\chi(x_{k})\geq\epsilon$ for all $k\in\{0,\dots,j_{\epsilon}-1\}$.
Denoting by $Nf_{j_{\epsilon}}$ the number of function evaluations required by Algorithm 1 up to iteration $j_{\epsilon}$, then $Nf_{j_{\epsilon}}\leq\operatorname{\mathcal{O}}{(n^{2}\epsilon^{-2})}$. In particular,
$Nf_{j_{\epsilon}}\leq 2n\left\lfloor\frac{nc_{3}^{2}(\Phi_{0}-f_{\min})}{c_{1}}\epsilon^{-2}\right\rfloor+\left\lfloor\frac{nc_{3}^{2}(f_{0}-f_{\min})}{\gamma c^{2}}\frac{\delta^{2}}{(1-\delta)^{2}}\epsilon^{-2}\right\rfloor,$
where $c_{1}$ and $c_{3}$ are given in (19) and (27), respectively.
Proof.
First, let us partition the set of iteration indices $\{0,\dots,j_{\epsilon}-1\}$ into $\mathcal{S}_{j_{\epsilon}}$ and $\mathcal{U}_{j_{\epsilon}}$ such that
$k\in\mathcal{S}_{j_{\epsilon}}\,\Leftrightarrow\,x_{k}\not=x_{k-1},\quad k\in\mathcal{U}_{j_{\epsilon}}\,\Leftrightarrow\,x_{k}=x_{k-1},\quad\mathcal{S}_{j_{\epsilon}}\cup\mathcal{U}_{j_{\epsilon}}=\{0,\dots,j_{\epsilon}-1\},$
that is, $\mathcal{S}_{j_{\epsilon}}$ and $\mathcal{U}_{j_{\epsilon}}$ contain the successful and unsuccessful iterations up to $j_{\epsilon}-1$, respectively.
When the algorithm evaluates a new point, the latter can either succeed to decrease the objective function or fail to do so. Let us then define $Nf_{j_{\epsilon}}^{\mathcal{S}}$ as the total number of function evaluations related to points which succeed to decrease the objective function up to iteration $j_{\epsilon}$. Note that, at each iteration, the maximum number of function evaluations related to points which fail to decrease the objective function is $2n$ (and it can be equal to $2n$ only when $T_{k}=\mathbb{R}^{n}$). So, we can write
$Nf_{j_{\epsilon}}\leq 2nj_{\epsilon}+Nf_{j_{\epsilon}}^{\mathcal{S}}\leq 2n\biggl{\lfloor}\frac{nc_{3}^{2}(\Phi_{0}-f_{\min})}{c_{1}}\epsilon^{-2}\biggr{\rfloor}+Nf_{j_{\epsilon}}^{\mathcal{S}},$ (30)
where the last inequality follows from Theorem 6. Now, let us consider any iteration $k<j_{\epsilon}$ and any index $i\in\{1,\ldots,n\}$ such that the line search succeeds to produce a decrease in the objective function. For each $\alpha$ used in the extrapolation phase of the line search, we have that
$f(y_{k}^{i}+\alpha d_{k}^{i})-f(y_{k}^{i}+(\alpha/\delta)d_{k}^{i})\geq\gamma\Biggl{(}\frac{1-\delta}{\delta}\Biggr{)}^{2}\alpha^{2}\geq\gamma\Biggl{(}\frac{1-\delta}{\delta}\Biggr{)}^{2}c^{2}\Delta_{k}^{2}.$ (31)
Let us define the index $m(k)$ as follows:
  • if $\mathcal{U}_{j_{\epsilon}}\cap\{0,\ldots,k-1\}\not=\emptyset$, then $m(k)$ is the largest index of $\mathcal{U}_{j_{\epsilon}}\cap\{0,\ldots,k-1\}$;
  • otherwise, $m(k)=0$.
Note that, by definition, $m(k)$ is the last unsuccessful iteration before iteration $k$, i.e., all the iterations from $m(k)+1$ to $k$ are successful iterations. Lemma 1 guarantees that $\Delta_{k}\geq\Delta_{m(k)+1}$ for all $k\in\mathcal{S}_{j_{\epsilon}}$. Hence, from (31), it follows that
$f(y_{k}^{i}+\alpha d_{k}^{i})-f(y_{k}^{i}+(\alpha/\delta)d_{k}^{i})\geq\gamma\Biggl{(}\frac{1-\delta}{\delta}\Biggr{)}^{2}\alpha^{2}\geq\gamma\Biggl{(}\frac{1-\delta}{\delta}\Biggr{)}^{2}c^{2}\Delta_{m(k)+1}^{2}.$
From Theorem 2, we have that $\Delta_{m(k)+1}\geq\chi(x_{m(k)})/(\sqrt{n}c_{3})$. Since $\chi(x_{k})\geq\epsilon$ for all $k\in\{0,\dots,j_{\epsilon}-1\}$, we can write
$f(y_{k}^{i}+\alpha d_{k}^{i})-f(y_{k}^{i}+(\alpha/\delta)d_{k}^{i})\geq\gamma\Biggl{(}\frac{1-\delta}{\delta}\Biggr{)}^{2}c^{2}\frac{\epsilon^{2}}{nc_{3}^{2}}.$
Then, recalling Assumption 1 and summing up the above relation over all function evaluations producing an objective decrease, we obtain
$f_{0}-f_{\min}\geq Nf_{j_{\epsilon}}^{\mathcal{S}}\gamma\Biggl{(}\frac{1-\delta}{\delta}\Biggr{)}^{2}c^{2}\frac{\epsilon^{2}}{nc_{3}^{2}},$
that is,
$Nf_{j_{\epsilon}}^{\mathcal{S}}\leq\left\lfloor\frac{nc_{3}^{2}(f_{0}-f_{\min})}{\gamma c^{2}}\frac{\delta^{2}}{(1-\delta)^{2}}\epsilon^{-2}\right\rfloor.$
The desired results hence follows from (30). ∎
4 Finite active-set identification
In this section, we show that Algorithm 1 identifies the components of the final solution lying on the lower or the upper bounds (the so called active set) in a finite number of iterations.
First, let us give an equivalent definition of stationarity for problem (1), which will be useful in our analysis.
Definition 3.
A point $x^{*}\in[l,u]$ is a stationary point of problem (1) (i.e., $\chi(x^{*})=0$) if and only if, for all $i\in\{1,\ldots,n\}$, we have that
$\nabla_{i}f(x^{*})\begin{cases}\geq 0\quad\text{if }x^{*}_{i}=l_{i},\\ =0\quad\text{if }l_{i}<x^{*}_{i}<u_{i},\\ \leq 0\quad\text{if }x^{*}_{i}=u_{i}.\end{cases}$
Now, let us recall the definition of strict complementarity and non-degenerate solutions.
Definition 4.
Given a stationary point $x^{*}$ of problem (1), we say that a component $x^{*}_{i}$ satisfies the strict complementarity condition if $x^{*}_{i}\in\{l_{i},u_{i}\}$ and $\nabla_{i}f(x^{*})\neq 0$. If the strict complementarity condition is satisfied by all components $x^{*}_{i}$, we say that $x^{*}$ is non-degenerate.
In particular, we define $\mathscr{Z}(x^{*})$ as the active set for a stationary point $x^{*}$, that is, the index set for the active components of $x^{*}$. We also define $\mathscr{Z}^{+}(x^{*})$ as the index set for those components satisfying the strict complementarity condition. Namely,
$\mathscr{Z}(x^{*})=\{i\colon x^{*}_{i}=l_{i}\}\cup\{i\colon x^{*}_{i}=u_{i}\}\qquad\text{and}\qquad\mathscr{Z}^{+}(x^{*})=\mathscr{Z}(x^{*})\cap\{i\colon\nabla_{i}f(x^{*})\neq 0\}.$
Furthermore, for any stationary point $x^{*}$ such that $\mathscr{Z}^{+}(x^{*})\neq\emptyset$, let us define
$\zeta(x^{*})=\min_{i\in\mathscr{Z}^{+}(x^{*})}\lvert\nabla_{i}f(x^{*})\rvert.$ (32)
We see that $\zeta(x^{*})$ is a measure of the minimum amount of strict complementarity among the variable in $\mathscr{Z}^{+}(x^{*})$. This quantity will be used to define a neighborhood of $x^{*}$ where the active components are correctly identified, following a similar approach as in [13, 32].
Before diving into the main theorem of this section, we need some preliminary results following from the Lipschitz continuity of $\nabla f$. Recalling Defintion 1, using standard arguments (see, e.g., [31]) one can prove that, for all $x\in[l,u]$, we have
$\lvert f(x+se_{i})-f(x)-s\nabla_{i}f(x)\rvert\leq\frac{L_{i}}{2}s^{2}\quad\forall s\in\mathbb{R}\colon x+se_{i}\in[l,u],\quad i=1,\ldots,n.$
Hence, for all $x\in[l,u]$, we have
$f(x+se_{i})$ $\leq f(x)+s\nabla_{i}f(x)+\frac{L_{i}}{2}s^{2}\quad\forall s\in\mathbb{R}\colon x+se_{i}\in[l,u],\quad i=1,\ldots,n,$ (33)
$f(x+se_{i})$ $\geq f(x)+s\nabla_{i}f(x)-\frac{L_{i}}{2}s^{2}\quad\forall s\in\mathbb{R}\colon x+se_{i}\in[l,u],\quad i=1,\ldots,n.$ (34)
The two following results provide bounds for the objective function when exploring any coordinate direction.
Proposition 5.
Given $x\in[l,u]$, $\gamma\geq 0$ and $i\in\{1,\ldots,n\}$, then
$f(x-s\operatorname{sign}(\nabla_{i}f(x))e_{i})\leq f(x)-\gamma s^{2}$
for all $0\leq s\leq 2\dfrac{\lvert\nabla_{i}f(x)\rvert}{L_{i}+2\gamma}$ such that $x-s\operatorname{sign}(\nabla_{i}f(x))e_{i}\in[l,u]$.
Proof.
From (33), we can write
$f(x+se_{i})\leq f(x)+s\biggl{(}\nabla_{i}f(x)+\frac{L_{i}}{2}s\biggr{)}\quad\forall s\in\mathbb{R}\colon x+se_{i}\in[l,u].$
The right-hand side of the above inequality is less than or equal to $f(x)-\gamma s^{2}$ if
$s\biggl{(}\nabla_{i}f(x)+\frac{L_{i}}{2}s\biggr{)}\leq-\gamma s^{2}.$
If $\nabla_{i}f(x)\neq 0$, solving with respect to $s$ we obtain
$-\frac{2\nabla_{i}f(x)}{L_{i}+2\gamma}\leq s\leq 0\quad$ if $\text{if }\nabla_{i}f(x)>0,$
$0\leq s\leq-\frac{2\nabla_{i}f(x)}{L_{i}+2\gamma}\quad$ if $\text{if }\nabla_{i}f(x)<0,$
leading to the desired result. ∎
Proposition 6.
Given $x\in[l,u]$ and $i\in\{1,\ldots,n\}$, then
$f(x+s\operatorname{sign}(\nabla_{i}f(x))e_{i})\geq f(x)$
for all $0\leq s\leq\dfrac{2\lvert\nabla_{i}f(x)\rvert}{L_{i}}$ such that $x+s\operatorname{sign}(\nabla_{i}f(x))e_{i}\in[l,u]$.
Proof.
From (34), we can write
$f(x+se_{i})\geq f(x)+s\biggl{(}\nabla_{i}f(x)-\frac{L_{i}}{2}s\biggr{)}\quad\forall s\in\mathbb{R}\colon x+se_{i}\in[l,u].$
The right-hand side of the above inequality is greater than or equal to $f(x)$ if
$s\biggl{(}\nabla_{i}f(x)-\frac{L_{i}}{2}s\biggr{)}\geq 0.$
If $\nabla_{i}f(x)\neq 0$, solving with respect to $s$ we obtain
$0\leq s\leq\frac{2\nabla_{i}f(x)}{L_{i}}\quad$ if $\text{if }\nabla_{i}f(x)>0,$
$\frac{2\nabla_{i}f(x)}{L_{i}}\leq s\leq 0\quad$ if $\text{if }\nabla_{i}f(x)<0,$
leading to the desired result. ∎
The next proposition shows that, when $\nu^{i}_{k}$ is sufficiently small at a given iteration, Algorithm 1 cannot move along an ascent direction.
Proposition 7.
Consider an iteration $k$ of Algorithm 1. If $\nu^{i}_{k}\leq 2\lvert\nabla_{i}f(y_{k}^{i})\rvert/L_{i}$ for an index $i\in\{1,\ldots,n\}$, then
$\alpha_{k}^{i}>0\;\Rightarrow\;d_{k}^{i}=-\operatorname{sign}(\nabla_{i}f(y_{k}^{i}))e_{i}.$
Proof.
Using Proposition 6, for all $\alpha\leq\nu^{i}_{k}$ and $\gamma>0$ we have
$f(y_{k}^{i}+\alpha\operatorname{sign}(\nabla_{i}f(y_{k}^{i}))e_{i})\geq f(y_{k}^{i})>f(y_{k}^{i})-\gamma\alpha^{2}.$
That is the line search in Algorithm 2 fails when using the direction $\operatorname{sign}(\nabla_{i}f(y_{k}^{i}))e_{i}$ with any stepsize $0<\alpha\leq\nu^{i}_{k}$. So, if the line search returns $\alpha_{k}^{i}>0$, necessarily $d_{k}^{i}=-\operatorname{sign}(\nabla_{i}f(y_{k}^{i}))e_{i}$. ∎
Now, we are ready to state the main result of this section, establishing finite active-set idenfitication of Algorithm 1.
Theorem 8.
Let $\{x_{k}\}$ be the sequence of points produced by Algorithm 1 and let $x^{*}$ be a limit point of $\{x_{k}\}$, i.e., there exists an infinite subsequence $\{x_{k}\}_{K}\to x^{*}$. Then,
  • (i)
    ${\lim_{k\to\infty,\,k\in K}x_{k+1}=x^{*}}$;
  • (ii)
    an iteration $\bar{k}\in K$ exists such that, for all $k\geq\bar{k}$, $k\in K$, we have that $(x_{k+1})_{i}=x^{*}_{i}$ for all $i\in\mathscr{Z}^{+}(x^{*})$.
Proof.
Since $y_{k}^{1}=x_{k}$ and $y_{k}^{i+1}=x_{k}+\sum_{j=1}^{i}\alpha_{k}^{j}d_{k}^{j}$, from Proposition 4 and the fact that $\lVert d_{k}^{i}\rVert=1$, $i=1,\ldots,n+1$, we have
$\lim_{\begin{subarray}{c}k\to\infty\\ k\in K\end{subarray}}y_{k}^{i}=\lim_{\begin{subarray}{c}k\to\infty\\ k\in K\end{subarray}}x_{k}=x^{*},\quad i=1,\ldots,n+1.$ (35)
Since $x_{k+1}=y_{k}^{n+1}$, then point (i) follows.
To show point (ii), assume that $\mathscr{Z}^{+}(x^{*})\neq\emptyset$. Let $\bar{k}\in K$ be the first iteration such that the two following relations hold for all $k\geq\bar{k}$, $k\in K$:
$\lVert y_{k}^{i}-x^{*}\rVert\leq\min\biggl{\{}\frac{1}{L},\frac{2}{2L+L^{\text{max}}+2\gamma}\biggr{\}}\zeta(x^{*}),\quad i=1,\ldots,n,$ (36a)
$\lVert y_{k}^{i}-x^{*}\rVert+\frac{L^{\text{max}}}{2L}\max_{j=1,\ldots,n}\tilde{\alpha}^{j}_{k}\leq\frac{\zeta(x^{*})}{L},\quad i=1,\ldots,n.$ (36b)
Note that (35) and Proposition 4 imply the existence of $\bar{k}\in K$ such that (36) holds for all $k\geq\bar{k}$, $k\in K$.
Consider an index $i\in\mathscr{Z}^{+}(x^{*})$ and an iteration $k\geq\bar{k}$, $k\in K$. To prove that $(x_{k+1})_{i}=x^{*}_{i}$, we have to show that $(y^{i+1}_{k})_{i}=x^{*}_{i}$ since, from the instructions of the algorithm, $(x_{k+1})_{i}=(y^{i+1}_{k})_{i}$. Without loss of generality, assume that $x^{*}_{i}=l_{i}$ (the proof for the case $x^{*}_{i}=u_{i}$ is identical, except for minor changes). So, we have to show that
$(y^{i+1}_{k})_{i}=l_{i}.$ (37)
Preliminarily, we want to prove that
$\lvert z_{i}-l_{i}\rvert\leq\frac{2\nabla_{i}f(z)}{L_{i}+2\gamma}\quad\forall z\text{ such that }\|z-x^{*}\|\leq\|y_{k}^{i}-x^{*}\|,$ (38)
$\nu^{i}_{k}\leq\frac{2\nabla_{i}f(y_{k}^{i})}{L_{i}}.$ (39)
According to Definitions 3 and 4, we have that $\nabla_{i}f(x^{*})>0$ and, from the definition of $\zeta(x^{*})$ given in (32), it follows that
$0<\zeta(x^{*})\leq\nabla_{i}f(x^{*}).$ (40)
Consider any $z\in\mathbb{R}^{n}$ such that $\|z-x^{*}\|\leq\|y_{k}^{i}-x^{*}\|$. Using the Lipschitz continuity of $\nabla f$, we have
$\nabla_{i}f(x^{*})-\nabla_{i}f(z)\leq\lVert\nabla f(z)-\nabla f(x^{*})\rVert\leq L\lVert z-x^{*}\rVert.$ (41)
Moreover, from (36a) we can write
$\|z-x^{*}\|\leq\|y_{k}^{i}-x^{*}\|\leq\frac{2\zeta(x^{*})}{2L+L^{\text{max}}+2\gamma}.$
Multiplying the first and last terms above by $(2L+L^{\text{max}}+2\gamma)/(L^{\text{max}}+2\gamma)$, we have
$\Bigl{(}\frac{2L}{L^{\text{max}}+2\gamma}+1\Bigr{)}\lVert z-x^{*}\rVert\leq\frac{2\zeta(x^{*})}{L^{\text{max}}+2\gamma},$
that is,
$\lVert z-x^{*}\rVert\leq\bigl{(}\zeta(x^{*})-L\lVert z-x^{*}\rVert\bigr{)}\frac{2}{L^{\text{max}}+2\gamma}.$ (42)
Since, from (40) and (41), we have
$\zeta(x^{*})\leq\nabla_{i}f(z)+L\lVert z-x^{*}\rVert,$ (43)
then, using (42), we obtain
$\lVert z-x^{*}\rVert\leq\frac{2\nabla_{i}f(z)}{L^{\text{max}}+2\gamma}.$
Taking into account that $L_{i}\leq L^{\text{max}}$ and recalling that $x^{*}_{i}=l_{i}$, it follows that (38) holds. To prove (39), from (36b) and the definition of $\nu^{i}_{k}$ we can write
$\nu^{i}_{k}\leq\max_{j=1,\ldots,n}\tilde{\alpha}^{j}_{k}\leq\bigl{(}\zeta(x^{*})-L\lVert y^{i}_{k}-x^{*}\rVert\bigr{)}\frac{2}{L^{\text{max}}}.$
Using (43) with $z=y^{i}_{k}$ and the fact that $L_{i}\leq L^{\text{max}}$, we thus get (39).
In view of (39) and Proposition 7, it follows that
$d^{i}_{k}=-e_{i},$ (44)
that is, $y^{i+1}_{k}=y^{i}_{k}-\alpha^{i}_{k}e_{i}$. Using $z=y^{i}_{k}$ in (38), we also have
$\alpha_{\text{max}}=\lvert(y^{i}_{k})_{i}-l_{i}\rvert\leq\frac{2\nabla_{i}f(y^{i}_{k})}{L_{i}+2\gamma},$ (45)
where $\alpha_{\text{max}}$ is the largest feasible stepsize along the direction $d^{i}_{k}$ at $y^{i}_{k}$. So, if $\alpha_{\text{max}}=0$, then $\alpha_{k}^{i}=0$ and, using (44), we have $(y_{k}^{i+1})_{i}=(y_{k}^{i})_{i}=l_{i}$, thus proving (37). If $\alpha_{\text{max}}>0$, using Proposition 5, (44) and (45), it follows that a sufficient decrease of $f$ along $d^{i}_{k}$ is obtained with the first stepsize $\bar{\alpha}$ used in the line search, that is, the condition at line 4 of Algorithm 2 is satisfied. Now, consider the extrapolation phase in the line search procedure, that is, lines 12–15 of Algorithm 2. Recalling (44), each stepsize $\eta=\min\{\alpha/\delta,\alpha_{\text{max}}\}$ is such that $\alpha\leq\eta\leq(y^{i}_{k})_{i}-l_{i}$, that is,
$0\leq\eta-\alpha\leq(y^{i}_{k}+\alpha d^{i}_{k})_{i}-l_{i}.$
So, from (44) and the fact that $x^{*}_{i}=l_{i}$, it follows that $\|y^{i}_{k}+\alpha d^{i}_{k}-x^{*}\|\leq\|y^{i}_{k}-x^{*}\|$. So, we can apply (38) with $z=y^{i}_{k}+\alpha d^{i}_{k}$ and then we obtain
$0\leq\eta-\alpha\leq(y^{i}_{k}+\alpha d^{i}_{k})_{i}-l_{i}\leq\frac{2\nabla_{i}f(y^{i}_{k}+\alpha d^{i}_{k})}{L_{i}+2\gamma}$
for every stepsize $\eta$ used in the extrapolation. Then, using Proposition 5 with $x=y^{i}_{k}+\alpha d^{i}_{k}$, $s=\eta-\alpha$ and $d^{i}_{k}$ as in (44), it follows that
$f(y^{i}_{k}+\eta d^{i}_{k})\leq f(y^{i}_{k}+\alpha d^{i}_{k})-\gamma(\eta-\alpha)^{2}.$
Namely, a sufficient decrease of $f$ is obtained with all stepsizes used in the extrapolation and we quit when we get the largest feasible stepsize, meaning that $(y^{i+1}_{k})_{i}$ will be at the lower bound $l_{i}$ and thus proving (37). ∎
Note that Theorem 8 establishes finite identification for any limit point of $\{x_{k}\}$, thus not requiring the convergence of the whole sequence. Note also, in the proof of Theorem 8, the crucial role played by the extrapolation in the line search procedure. Loosely speaking, when we are sufficiently close to a stationary point, expanding the stepsize allows us to hit the lower or the upper bound, provided the strict complementarity condition holds. This guarantees to identify all the variables satisfying the strict complementarity after a finite number of iterations.
Now, let us point out a useful property for the limit points of $\{x_{k}\}$. To this aim, let us define $X^{\star}$ as the set of all limit points of $\{x_{k}\}$, i.e.,
$X^{\star}:=\{x\colon\exists K\subseteq\mathbb{N}\text{ such that }\lim_{k\to\infty,k\in K}x_{k}=x\}.$
The following result roughly states that, if for all $x\in X^{\star}$ there does not exists $i$ such that $x_{i}$ violates the strict complementarity, then either all $x_{i}$ lie on the same bound or they all are strictly feasible.
Proposition 8.
Let $\{x_{k}\}$ be the sequence of points produced by Algorithm 1 and consider an index $i\in\{1,\ldots,n\}$. Assume that there is no $x\in X^{\star}$ such that $i\in\mathscr{Z}(x)\setminus\mathscr{Z}^{+}(x)$. Then,
  • if there exists $x^{*}\in X^{\star}$ such that $i\in\mathscr{Z}^{+}(x^{\star})$, we have that $x_{i}=x^{*}_{i}$ for all $x\in X^{\star}$;
  • otherwise, $x_{i}\in(l_{i},u_{i})$ for all $x\in X^{\star}$.
Proof.
We limit ourselves to show only the first point since the second one can be obtained as a logical consequence. From Theorem 4 and recalling Ostrowski’s theorem [4], the set $X^{\star}$ of limit points of the sequence $\{x_{k}\}$ is a connected set. Now, let us consider any two points in $X^{\star}$, say $\bar{x}\in X^{\star}$ and $\bar{y}\in X^{\star}$, such that $\bar{x}\neq\bar{y}$, and $\bar{y}_{i}\in\{l_{i},u_{i}\}$. Since $X^{\star}$ is connected, there exists a continuous function $\rho:[a,b]\to\mathbb{R}^{n}$ such that $\rho(a)=\bar{x}$, $\rho(b)=\bar{y}$ and $\rho(t)\in X^{\star}$, i.e., $\rho(t)$ is stationary, for all $t\in[a,b]$. Let us assume, without loss of generality, that $\bar{y}_{i}=l_{i}$ (the proof for the case $\bar{y}_{i}=u_{i}$ is identical, except for minor changes). By contradiction, now assume that $\bar{x}_{i}>l_{i}$. Since $\rho(a)_{i}=\bar{x}_{i}>l_{i}$ and $\rho(b)_{i}=\bar{y}_{i}=l_{i}$, then there exists $\bar{t}\in(a,b]$ such that $\rho(t)_{i}>l_{i}$ for all $t\in[a,\bar{t})$ and $\rho(\bar{t})_{i}=l_{i}$. Furthermore, by the stationarity conditions given in Definition 3, we have that $\nabla_{i}f(\rho(t))=0$ for all $t\in[a,\bar{t})$ and $\nabla_{i}f(\rho(\bar{t}))>0$, where the last inequality follows from the stated hypothesis. Then, by continuity of $\nabla f$, a scalar $\hat{t}\in(a,\bar{t})$ must exist such that $\nabla_{i}f(\rho(t))>0$ for all $t\in(\hat{t},\bar{t}]$. This is a contradiction since $\nabla_{i}f(\rho(t))=0$ for all $t\in(\hat{t},\bar{t})$. ∎
Applying the above proposition for all indices $i\in\{1,\ldots,n\}$, the following result immediately follows, enforcing the finite active-set identification property established in Theorem 8 when all the limit points of $\{x_{k}\}$ are non-degenerate.
Corollary 1.
Let $\{x_{k}\}$ be the sequence of points produced by Algorithm 1 and assume that every $x\in X^{\star}$ is non-degenerate. Then, for any pair $x^{\prime},x^{\prime\prime}\in X^{\star}$, we have $\mathscr{Z}(x^{\prime})=\mathscr{Z}(x^{\prime\prime})$ and $x^{\prime}_{i}=x^{\prime\prime}_{i}$ for all $i\in\mathscr{Z}(x^{\prime})$.
5 Conclusions
In this paper, we have analyzed a derivative-free line search method for bound-constrained problems where the objective function has a Lipschitz continuous gradient. For this algorithm, we have first provided complexity results. In more detail, given a prespecified threshold $\epsilon>0$, we have shown that the criticality measure $\chi(x_{k})$ (which vanishes at stationary points) falls below $\epsilon$ after at most $\operatorname{\mathcal{O}}{(n\epsilon^{-2})}$ iterations, requiring at most $\operatorname{\mathcal{O}}{(n^{2}\epsilon^{-2})}$ function evaluations. These bounds match those obtained for (deterministic) direct-search [19] and model-based [22] methods. Additionally, we have established an upper bound of $\operatorname{\mathcal{O}}{(n^{2}\epsilon^{-2})}$ on the total number of iterations where $\chi(x_{k})\geq\epsilon$. The latter result is obtained thanks to the extrapolation strategy used in the proposed line search, allowing us to upper bound $\chi(x_{k})$ on both successful and unsuccessful iterations.
In the last part of the paper, we have considered the active-set identification property of the proposed method, i.e., the capability to detect the variables lying at the lower or the upper bound in the final solutions. In this respect, we have shown that, in a finite number of iterations, the algorithm identifies the active constraints satisfying the strict complementarity condition. Also this property is obtained by exploiting the extrapolation used in the proposed line search, allowing the stepsize to expand, when we are in a neighborhood of a stationary point, until we hit the boundary of the feasible set.
Finally, some topics for future research can be envisaged. In particular, under convexity assumptions, the worst-case complexity of the algorithm might be tightened, in order to match the results given in [15], and a bound on the maximum number of iterations required to identify the active constraints might be given. We wish to report more results in future works.